Pls. help!!! I can't understand the question.
On top of the building, a cylindrical tank is 22 ft high and with a radius of 6 ft. is filled with water. If water has a mass of 62.4 lbs/cubic ft., what is the mass of the water in the tank when it is full?
I need a formula and a pattern to understand this. Pls. help me answer this question. Thanks.
: Pls. help!!! I can't understand the question.
On top of the building, a cylindrical tank is 22 ft high and with a radius of 6 ft. is filled with water. If water has a mass of 62.4 lbs/cubic ft., what is the mass of the water in the tank when it is full?
I need a formula and a pattern to understand this. Pls. help me answer this question. Thanks.

Answer by nerdybill(1129)

On top of the building, a cylindrical tank is 22 ft high and with a radius of 6 ft. is filled with water. If water has a mass of 62.4 lbs/cubic ft., what is the mass of the water in the tank when it is full?
.
First, you must determine the "volume" of the cylindrical tank.
Volume for a cylinder is:
volume = (pi)r^2
where
pi is 3.14
r is radius
.
So, for your tank:
Volume = (31.4)(6)^2 = (31.4)(36) = 1130.4 cubic ft
.
The problem gave you the "mass" of water as 62.4 lbs/cubic ft -- simply multiply this by the volume and you'll get your answer:
62.4 lbs/cubic ft * 1130.4 cubic ft = 70536.96 lbs